OpenStudy (anonymous):

need to find the solution to this problem?? can anyone help me? will give medal

OpenStudy (anonymous):

\[4 \left[ x + 3 \right] - 8 = 8 \]

OpenStudy (anonymous):

first step is to try and get the x by its self

OpenStudy (anonymous):

so to do that get rid of anything that is being added or subtracted to X

OpenStudy (anonymous):

that would be the 1st step in getting x all by its self...so what would u do...

OpenStudy (anonymous):

The first step is to Simplify or Merge the same terms.

OpenStudy (mathmale):

Peyton! by [x+3] do you mean "absolute value of x+3" or something else?

OpenStudy (mathmale):

If it were absolute value, you could type |x+3| instead.

OpenStudy (mathmale):

Peyton! Need to hear from you. All of us do.

OpenStudy (anonymous):

its absolute value

OpenStudy (mathmale):

In that case, Peyton, add 8 to both sides of the given equation, and combine terms on the right side. GirlByte: Would you plese check your answer in the original equation?

OpenStudy (anonymous):

honestly neither of those are my choices guys

OpenStudy (anonymous):

so close u just forgot to add the 8 to the other side

OpenStudy (anonymous):

of yeah the answer is 1

OpenStudy (mathmale):

Peyton: I wasn't giving you a choice. I was asking you to write down the original relationship and then add 8 manually to both sides of the equation. Perhaps Paul can walk you through the rest of the solution; I for one need to hit the sack.

OpenStudy (mathmale):

GirlByte: you say you've substituted x=5 into the original equation and get the answer 1? What we're supposed to do here is to substitute x=5 into the original equation and ask ourselves wehther the resulting relationship is true or not.

OpenStudy (anonymous):

nah i gotta go too good luck girl! :)

OpenStudy (mathmale):

Here's what I would do: 4|x+3|-8=8 I asked you to add 8 to both sides of the equation. Then 4|x+3|-8+8=8+8 And then 4|x+3|=16 Divide both sides by 4 to get |x+3|=4. This has TWO solutions. Could some of you out there brainstorm on how to find them? Note: 5 is NOT one of the answers.

OpenStudy (anonymous):

heres are my choices